OpenStudy (emmyma):

*help please! ^^ will give a medal and fan!!!* [(x+3)/x] times [(x+4)/(x^2+7x+12)]

OpenStudy (emmyma):

\[\frac{ x+3 }{ x} * \frac{ x+4 }{ x^{2}+7x+12 }\] Thank you so much for helping!!

jimthompson5910 (jim_thompson5910):

What do you get when you factor x^2 + 7x + 12

OpenStudy (emmyma):

factoring x^2 + 7x + 12 is -4, and -3.

jimthompson5910 (jim_thompson5910):

two numbers that multiply to 12 and add to 7 are 4 and 3 so that means x^2 + 7x + 12 factors to (x+4)(x+3)

jimthompson5910 (jim_thompson5910):

\[\large \frac{ x+3 }{ x} * \frac{ x+4 }{ x^{2}+7x+12 }\] \[\large \frac{ x+3 }{ x} * \frac{ x+4 }{ (x+3)(x+4) }\] which terms cancel?

OpenStudy (emmyma):

i'm not really getting this, could you try to explain it further please? thank you:)

jimthompson5910 (jim_thompson5910):

Do you see how x^2 + 7x + 12 factors to (x+4)(x+3) ?

OpenStudy (emmyma):

yes, i understand that!

jimthompson5910 (jim_thompson5910):

so we factor and then we cancel like so... \[\large \frac{ x+3 }{ x} * \frac{ x+4 }{ x^{2}+7x+12 }\] \[\large \frac{ x+3 }{ x} * \frac{ x+4 }{ (x+3)(x+4) }\] \[\large \frac{ x+3 }{ x} * \frac{ \cancel{x+4} }{ (x+3)\cancel{(x+4)} }\] \[\large \frac{ x+3 }{ x} * \frac{ 1 }{ x+3 }\] \[\large \frac{ \cancel{x+3} }{ x} * \frac{ 1 }{ \cancel{x+3} }\] \[\large \frac{ 1 }{ x} * \frac{ 1 }{ 1 }\] \[\large \frac{ 1 }{ x} * 1\] \[\large \frac{ 1 }{ x}\] ------------------------------------------------------------------ This means \[\large \frac{ x+3 }{ x} * \frac{ x+4 }{ x^{2}+7x+12 }\] simplifies to \[\large \frac{ 1 }{ x}\]

jimthompson5910 (jim_thompson5910):

The cancellations occur because when you divide any number by itself (excluding 0) you get 1 examples 2/2 = 1 7/7 = 1 22/22 = 1 in general x/x = 1 where x is not zero

jimthompson5910 (jim_thompson5910):

and the 1s go away because 1*x = x examples 1*9 = 9 1*17 = 17 etc etc

OpenStudy (emmyma):

I'm still not quite understanding where you are going with this? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you see the steps I listed out?

jimthompson5910 (jim_thompson5910):

that went from \[\large \frac{ x+3 }{ x} * \frac{ x+4 }{ x^{2}+7x+12 }\] to \[\large \frac{ 1 }{ x}\]

OpenStudy (emmyma):

yes, how would i continue to solve this problem?

jimthompson5910 (jim_thompson5910):

once you get to \[\large \frac{ 1 }{ x}\] you are done

jimthompson5910 (jim_thompson5910):

because it's fully simplified

OpenStudy (emmyma):

ahh, i see!! thank you so much for explaining it so thoroughly. :)

jimthompson5910 (jim_thompson5910):

you're welcome